OpenStudy (anonymous):

Solve: 4x − 10 + 16 = 22 4 12 7 −4

OpenStudy (kymber):

Did you try to solve it? Where are you getting stuck?

OpenStudy (anonymous):

-4?

jimthompson5910 (jim_thompson5910):

no, -4 is not the answer

OpenStudy (anonymous):

Answer is 4 It is very simple 4x-6=22 4x=22-6 4x=16 x=16/4 x=4

OpenStudy (kymber):

Do you know what you did wrong?

OpenStudy (anonymous):

This is a simple linear equation in one variable

OpenStudy (anonymous):

Ohhh! I Feel Dumb!

jimthompson5910 (jim_thompson5910):

just keep practicing and you'll be fine, and you're far from dumb so don't let anyone tell you that

OpenStudy (anonymous):

you are not dumb i still remember my old math classes I had to practise a lot to get the hang of it